# XAI-CF – Examining the Role of Explainable Artificial Intelligence in Cyber Forensics

### v2

@@ 6.1.1 Summary
The adoption of statistical methods in the field of CF is driven by various compelling reasons. These methods are instrumental in handling missing information and reducing errors, creating a robust foundation for decision-making in legal contexts. Moreover, statistical techniques contribute to the graphical representation of legal evidence, facilitating the visualization and interpretation of intricate relationships within the data. The integration of scenarios within Bayesian networks is a notable strength of statistical approaches, offering a comprehensive modeling strategy for understanding complex legal cases. In court settings, the use of statistical information, such as probabilities and likelihood ratios, becomes crucial for presenting and evaluating evidence, especially in cases involving complex data like DNA profiling. Furthermore, statistical methods enable pattern recognition, aiding in the identification of recurrent features in legal evidence. The complexity of legal cases is effectively managed through interpretable statistical models, enhancing transparency and ensuring rigorous scrutiny of AI decision-making processes. The practicality and accessibility of statistical models make them valuable tools for non-specialists, such as legal professionals, to interpret and utilize in the decision-making process. By minimizing the risk of miscarriages of justice and providing a systematic, evidence-based approach to evaluating legal evidence, XAI-CF methods offer indispensable support for informed and just decision-making in the legal domain.
@@ 6.2 Machine Learning (ML)
Machine Learning (ML) is a major and core part of an AI system. ML gives intelligence to machines and makes them partially think on their own. There are three basic approaches to ML, supervised, unsupervised, and reinforced learning. Here we are not going to go into the details of these approaches. For an introduction and a detailed discussion on ML, interested readers are referred to (Introduction-ML-2020, ). This section presents a review and discussion of some of the research works that explain AI systems in CF using ML.
@@
Khan et al. (khan2018eliciting, ) present an associative rule mining technique to analyze security event logs, which can be used during forensics investigation of security events. Firstly, event log entries are converted into an object-based model, and then associative rules are extracted dynamically from this model. Secondly, The quality of these rules is improved by filtering insignificant rules. This helps in forming and validating the sequence of events. For forensics investigation, these events can be further examined by a forensics expert. The examination can also be automated by using an action-based domain model. The paper highlights that vulnerability assessment and security configuration activities heavily depend on expert knowledge. Incorporating ML, especially XAI models in this scenario, could aid in automating certain aspects, making the expertise more accessible to non-professionals.
@@
Aditya et al. (aditya2018enabling, ) present an ATF (Adversary Testing Framework) to test the security robustness of black box DNN (Deep Neural Networks) models in the context of CF. The purpose of developing the ATF is to develop trust in a DNN model. The ATF probes a DNN model’s defenses and tries to figure out the workings of the model by calculating the influence of the features on the prediction. The Integrated Gradient (sundararajan2017axiomatic, ) technique and a surrogate model are used for this purpose. This helps the framework in crafting the inputs to the DNN model for testing its security robustness.
@@
Afzaliseresht et al. (afzaliseresht2019explainable, ) proposed an XAI model for analyzing security event logs, which can be used during forensics investigation of security events. An apriori algorithm (agrawal1994fast, ) is used to mine temporal patterns to automatically discover sequential events from a log file, a common application of ML. Storytelling (wu2013internet, ) is used to present this sequence of events as a knowledge model. This knowledge model highlights the semantic and inferred information from log files in a human-readable format. This provides a contextualized interpretation of security events which can help explain the events and hence reduces the efforts of humans to interpret events.
@@
Costantini et al. (costantini2020assessing, ) propose a formal framework for assessing the information quality of IoT devices. During forensic analysis of IoT devices using XAI, this framework provides more transparency to the reasoning process during evidence collection for investigation purposes. The authors develop a three-part analytical process for assessing the information quality. Firstly, the preservation of the integrity of collected information. Secondly, the trustworthiness of events is verified by the sources of pieces of evidence. Thirdly, the quality of the discussion about pieces of evidence depends on the competencies of the agents, argumentation abilities, etc. The paper highlights challenges in handling a vast number and extreme variety of IoT items, often lacking physical interfaces, pointing to the need for advanced computational methods. In this work, ML is applied, especially within the XAI-CF framework, which has contributed to handling the complexity and diversity of IoT data, and provided tools for automated analysis and interpretation.
@@
Bolle et al. (bolle2020role, ) presents a framework based on scientific interpretation principles to evaluate AI forensics analysis systems. This helps in building trust in such systems. The framework comprises three levels/steps, performance evaluation, understandability and transparency evaluation, and forensic evaluation. For the final forensic conclusion to be reliable and trustworthy each of these steps should be formalized. The authors also discussed challenges of measuring understandability, such as subjectivity, data protection, and the risk of making a system simpler to understand but missing an important aspect. The mention of undetected errors or bias resulting in wrong decisions in the paper (bolle2020role, ) emphasizes the need for transparency and interpretability in automated systems. In the context of forensic analysis, especially when utilizing ML, these techniques become crucial for ensuring that the decision-making process is understandable and reliable. Also, the reference to decisions violating fundamental human rights due to outputs that are not well understood highlights the ethical implications of automated systems. XAI-CF focuses on making AI systems interpretable and explainable, addressing concerns related to accountability and human rights.
@@
Mahajan et al. (mahajan2021explainable, ) evaluate different AI systems for toxic comment classification. During forensic analysis of social media content, there is a need to check the toxicity of the communication on social media platforms. During experiments, the authors noticed that the standard evaluation metrics, such as accuracy, precision, recall, etc, can be deceiving in evaluating AI systems. They instead use LIME, one of the XAI techniques for model interpretability to evaluate AI models. The results concluded that XAI techniques like LIME are important in selecting the best model and also increasing the user’s trust in the model. The paper underscores the significance of model interpretability techniques, specifically mentioning LIME, to understand the decisions made by ML models. This aligns with the core principles of XAI-CF, which emphasizes the interpretability of AI models, especially in sensitive domains like CF.
@@
Nguyen et al. (nguyen2021human, ) propose to use a human-in-the-loop (an expert forensic analyst) in an XAI system so that both of them complement each other to produce effective and efficient results. The forensic analyst compares the pieces of evidence and provides adjustments to the errors identified in the previous iteration. The input to the XAI model is modified according to these adjustments. This way with each iteration the system improves. In the end, the forensic analyst produces the final report with all the observable pieces of evidence. The proposed framework emphasizes the integration of security analysts or forensic investigators into the man-machine loop. The model is still under development and there are no experimental studies presented in the paper.
@@
Pethe et al. (pethe2022atle2fc, ) propose a model for explainable IoT forensics using ensemble learning. The use of an ensemble layer comprising various machine-learning models for classification demonstrates the incorporation of ML principles. Ensemble learning involves the combination of multiple models to enhance predictive performance, highlighting the relevance of the ML class. The ensemble layer consists of different ML models for producing the classification. The classification results from the ensemble layer are input to the explainable layer that consists of FPGrowth (Frequent Pattern Growth) and a deep learning classifier. FPGrowth is a recommender that assists in finding mutually dependent events. A deep learning model then categorizes the severity level of these events. The severity level in turn is used to fine-tune the classification and explainability models.
@@
Hall et al. (hall2022proof, ) evaluate different AI models using LIME. These models were trained to predict image and file content and file system metadata on a virtual hard disk image. These predictions were processed using LIME, an XAI tool that enhances the performance and investigative capabilities of AI’s model to assist CF professionals. The authors collected a sample of 23 virtual hard disk files containing a Windows system with a New Technology File System (NTFS) and a single user account. Different file types were created for forensics analysis. After classification, the results were input to LIME for explanation. The results indicate that LIME most of the time was able to explain the classification results but sometimes it failed, such as when feature interaction was involved. The use of LIME to evaluate different AI models for forensics analysis reinforces the connection of this research with XAI-CF.
@@
Jayakumar et al. (jayakumar2022visually, ) propose a method that enhances the interpretability of deepfake detection models. This method is employed to visually explain why a deepfake detection model classifies a video as a deepfake. The deep fake detection model could play a crucial role in the decision-making process of juries in digital forensic investigations. The authors use Anchors (ribeiro2018anchors, ) one of the XAI methods, emphasizing the integration of XAI techniques. This shows that XAI methods play a crucial role in providing interpretability and transparency to ML models, especially in contexts where justifications behind tool decisions are essential.
@@
De et al. (de2022interpretable, ) take a step towards explainable CF by proposing an interpretable federated transformer log learning model for threat detection in a cloud. The authors propose a model for threat detection in a cloud environment. This model is based on a local transformer-based threat detection model trained at each client within an organizational unit. Firstly, a local model is trained for each client using the local dataset of system logs. Secondly, the learned parameters from all the local models are fed to the federated learning server to generate a global federated learning model, which is shared with each client. This cycle is repeated many times to improve threat detection. Federated interpretability weights are computed by aggregating the local interpretability weights of each client. The interpretability module of each client highlights differences in normal and threat sequences that in turn help in the decision-making process. Additionally, it provides forensic investigators with actionable information to effectively analyze system activity and operation.
@@
Bouter et al. (bouter2023protoexplorer, ) present a system named Protoexplorer for visualizing and interpreting predictions of deepfake video data for forensics analysis. The paper mentions the importance of interpretable ML models in high-stakes settings, especially with the prevalence of complex deep-learning models. Also, the paper introduces prototype-based methods as a promising approach to make deep learning interpretable. Protoexplorer improves the explainability of deepfake video detection models by replacing the prototypes with alternative suggested prototypes. Protoexplorer provides a forensics analyst with meaningful visualizations and intuitive interactions with models. This helps the forensics analyst thoroughly evaluate and explain the models. The main action a forensics analyst can perform is zooming in on a part of the image and observing the contribution of the model to the prediction score. This step is called forensic deepfake analysis where a decision is made if the video is manipulated or not. This decision should be interpreted so that it can be presented in a court of law as a piece of trustworthy evidence. For interpretation, the forensics analyst explores and refines the model and provides an explanation and understanding of how the model aided in making the decision.
@@ 6.2.1 Summary
In the realm of high-stakes scenarios and complex technological landscapes, the inclusion of AI becomes imperative for a myriad of reasons. The deepening intricacies of Deep Learning Models demand a robust contextual understanding, and AI serves as the bedrock for developing, training and applying these models effectively. In the critical field of deepfake analysis, AI’s ability to discern patterns and anomalies is indispensable for creating algorithms capable of detecting and analyzing deceptive content. Additionally, AI takes center stage in the domains of threat detection and forensics, providing the computational prowess necessary for processing extensive datasets and uncovering patterns vital for identifying potential threats. The advent of XAI methods and interpretability techniques like LIME further solidifies AI’s role in CF, as it underpins the development of models that can be comprehensively understood and trusted by human users. Moreover, the utilization of DNN, ensemble learning, visualization, and human-in-the-loop showcase XAI’s versatility and indispensability in crafting effective solutions for multifaceted challenges. From addressing concerns related to undetected errors and bias to navigating the complexity and variety of IoT data, XAI emerges as a linchpin, contributing to ethical, interpretable, and powerful technological solutions across diverse domains in CF.
@@ 7 Challenges and Future Works
In this section, we discuss some of the challenges that need to be resolved for the promising development and use of a practical XAI-CF system. We also provide and recommend solutions to these challenges as part of future work.
7.0.1 Human Computer Interaction (HCI)
@@
Interaction with an XAI-CF system allows the user to ask follow-up questions/queries to better understand and evaluate the explanations. Another major reason for interacting with an XAI-CF system is to include a human-in-the-loop that can produce, valuable, practical, and structured explanations. To build such systems that are successful in practice we need to provide the user with a usable interactive XAI-CF system. Human-Computer Interaction (HCI) and XAI are upcoming research areas and very limited research (wolf2019explainability, ; shneiderman2020bridging, ) has been reported in this area and very few (bouter2023protoexplorer, ) on HCI and XAI-CF. As part of the CF research community, this is our responsibility to build the new XAI-CF systems interactive from the ground up. There are already several research works on interacting with AI systems (kulesza2012tell, ; kulesza2009fixing, ; lee2017human, ; stumpf2009interacting, ; yang2013learning, ). Recently developed ChatGPT (ray2023chatgpt, ; fui2023generative, ) the new AI language model has revolutionized the approach in AI to human-model-interaction. ChatGPT is being used in various domains, such as business, education, and healthcare, etc. The potentials of ChatGPT are unlimited, but it can be a double-edged sword (fui2023generative, ). For example, for businesses, it can create creative content such as ideas for advertisements, and on the other hand, it can create hallucinations and produce fake information. One of the future works is to explore the use of ChatGPT to improve HCI in XAI-CF systems and identify fake or false information for its successful use. Consideration of ethical standards and security measures should be paramount in the development process of interactive XAI-CF systems. A practical XAI-CF system is human-centered, because without human intervention and understanding CF produces no results. Therefore, HCI for such a system should be guided by human values and needs. One such challenge is socio-technical and is about interacting with an XAI-CF system by a person (e.g., law enforcement community) who has little knowledge of such systems. For example, retraining a model on new data is an advanced skill and requires setting up hyperparameters that influence the learning process. One solution is to transfer such knowledge to the law enforcement community or develop new HCI tools that facilitate and automate such actions with ease and without requiring advanced skills or knowledge. We recommend the latter option which makes the solution general and is human-centered.
7.0.2 Bias Management

### v3

@@ 6.1.1 Summary
The three papers focus on enhancing the interpretability and trustworthiness of forensic analyses through structured explanations. Vlek et al. [147] propose scenario-based explanations for Bayesian networks, providing probabilistic interpretations and evidential support to make the networks understandable to legal experts. Verheij et al. [146] explore hybrid models that connect arguments, scenarios, and probabilities, aiming to reduce reasoning errors and miscommunication between legal and forensic experts. Gopinath et al. [205] focus on IoT forensics, presenting quantitative metrics and detailed scenario reports that increase trust in digital evidence while requiring minimal effort from analysts. Collectively, these works highlight the importance of transparent, structured, and quantitative explanations to support reliable forensic decision-making.
@@ 6.2 Machine Learning (ML)
Machine Learning (ML) is a major and core part of an AI system. ML gives intelligence to machines and makes them partially think on their own. There are three basic approaches to ML, supervised, unsupervised, and reinforced learning. Here we are not going to go into the details of these approaches. For an introduction and a detailed discussion on ML, interested readers are referred to [210]. This section presents a review and discussion of some of the research works that explain AI systems in CF using ML.
@@
Khan et al. [204] present an associative rule mining technique to analyze security event logs, which can be used during forensics investigation of security events. Firstly, event log entries are converted into an object-based model, and then associative rules are extracted dynamically from this model. Secondly, The quality of these rules is improved by filtering insignificant rules. This helps in forming and validating the sequence of events. For forensics investigation, these events can be further examined by a forensics expert. The examination can also be automated by using an action-based domain model. The paper highlights that vulnerability assessment and security configuration activities heavily depend on expert knowledge. Incorporating ML, especially XAI models in this scenario, could aid in automating certain aspects, making the expertise more accessible to non-professionals.
@@
Aditya et al. [148] present an ATF (Adversary Testing Framework) to test the security robustness of black box DNN (Deep Neural Networks) models in the context of CF. The purpose of developing the ATF is to develop trust in a DNN model. The ATF probes a DNN model’s defenses and tries to figure out the workings of the model by calculating the influence of the features on the prediction. The Integrated Gradient [211] technique and a surrogate model are used for this purpose. This helps the framework in crafting the inputs to the DNN model for testing its security robustness.
@@
Afzaliseresht et al. [149] proposed an XAI model for analyzing security event logs, which can be used during forensics investigation of security events. An apriori algorithm [212] is used to mine temporal patterns to automatically discover sequential events from a log file, a common application of ML. Storytelling [213] is used to present this sequence of events as a knowledge model. This knowledge model highlights the semantic and inferred information from log files in a human-readable format. This provides a contextualized interpretation of security events which can help explain the events and hence reduces the efforts of humans to interpret events.
@@
Costantini et al. [150] propose a formal framework for assessing the information quality of IoT devices. During forensic analysis of IoT devices using XAI, this framework provides more transparency to the reasoning process during evidence collection for investigation purposes. The authors develop a three-part analytical process for assessing the information quality. Firstly, the preservation of the integrity of collected information. Secondly, the trustworthiness of events is verified by the sources of pieces of evidence. Thirdly, the quality of the discussion about pieces of evidence depends on the competencies of the agents, argumentation abilities, etc. The paper highlights challenges in handling a vast number and extreme variety of IoT items, often lacking physical interfaces, pointing to the need for advanced computational methods. In this work, ML is applied, especially within the XAI-CF framework, which has contributed to handling the complexity and diversity of IoT data, and provided tools for automated analysis and interpretation.
@@
Bolle et al. [151] presents a framework based on scientific interpretation principles to evaluate AI forensics analysis systems. This helps in building trust in such systems. The framework comprises three levels/steps, performance evaluation, understandability and transparency evaluation, and forensic evaluation. For the final forensic conclusion to be reliable and trustworthy each of these steps should be formalized. The authors also discussed challenges of measuring understandability, such as subjectivity, data protection, and the risk of making a system simpler to understand but missing an important aspect. The mention of undetected errors or bias resulting in wrong decisions in the paper [151] emphasizes the need for transparency and interpretability in automated systems. In the context of forensic analysis, especially when utilizing ML, these techniques become crucial for ensuring that the decision-making process is understandable and reliable. Also, the reference to decisions violating fundamental human rights due to outputs that are not well understood highlights the ethical implications of automated systems. XAI-CF focuses on making AI systems interpretable and explainable, addressing concerns related to accountability and human rights.
@@
Mahajan et al. [152] evaluate different AI systems for toxic comment classification. During forensic analysis of social media content, there is a need to check the toxicity of the communication on social media platforms. During experiments, the authors noticed that the standard evaluation metrics, such as accuracy, precision, recall, etc, can be deceiving in evaluating AI systems. They instead use LIME, one of the XAI techniques for model interpretability to evaluate AI models. The results concluded that XAI techniques like LIME are important in selecting the best model and also increasing the user’s trust in the model. The paper underscores the significance of model interpretability techniques, specifically mentioning LIME, to understand the decisions made by ML models. This aligns with the core principles of XAI-CF, which emphasizes the interpretability of AI models, especially in sensitive domains like CF.
@@
Nguyen et al. [159] propose to use a human-in-the-loop (an expert forensic analyst) in an XAI system so that both of them complement each other to produce effective and efficient results. The forensic analyst compares the pieces of evidence and provides adjustments to the errors identified in the previous iteration. The input to the XAI model is modified according to these adjustments. This way with each iteration the system improves. In the end, the forensic analyst produces the final report with all the observable pieces of evidence. The proposed framework emphasizes the integration of security analysts or forensic investigators into the man-machine loop. The model is still under development and there are no experimental studies presented in the paper.
@@
Pethe et al. [153] propose a model for explainable IoT forensics using ensemble learning. The use of an ensemble layer comprising various machine-learning models for classification demonstrates the incorporation of ML principles. Ensemble learning involves the combination of multiple models to enhance predictive performance, highlighting the relevance of the ML class. The ensemble layer consists of different ML models for producing the classification. The classification results from the ensemble layer are input to the explainable layer that consists of FPGrowth (Frequent Pattern Growth) and a deep learning classifier. FPGrowth is a recommender that assists in finding mutually dependent events. A deep learning model then categorizes the severity level of these events. The severity level in turn is used to fine-tune the classification and explainability models.
@@
Hall et al. [154] evaluate different AI models using LIME. These models were trained to predict image and file content and file system metadata on a virtual hard disk image. These predictions were processed using LIME, an XAI tool that enhances the performance and investigative capabilities of AI’s model to assist CF professionals. The authors collected a sample of 23 virtual hard disk files containing a Windows system with a New Technology File System (NTFS) and a single user account. Different file types were created for forensics analysis. After classification, the results were input to LIME for explanation. The results indicate that LIME most of the time was able to explain the classification results but sometimes it failed, such as when feature interaction was involved. The use of LIME to evaluate different AI models for forensics analysis reinforces the connection of this research with XAI-CF.
@@
Jayakumar et al. [155] propose a method that enhances the interpretability of deepfake detection models. This method is employed to visually explain why a deepfake detection model classifies a video as a deepfake. The deep fake detection model could play a crucial role in the decision-making process of juries in digital forensic investigations. The authors use Anchors [85] one of the XAI methods, emphasizing the integration of XAI techniques. This shows that XAI methods play a crucial role in providing interpretability and transparency to ML models, especially in contexts where justifications behind tool decisions are essential.
@@
De et al. [156] take a step towards explainable CF by proposing an interpretable federated transformer log learning model for threat detection in a cloud. The authors propose a model for threat detection in a cloud environment. This model is based on a local transformer-based threat detection model trained at each client within an organizational unit. Firstly, a local model is trained for each client using the local dataset of system logs. Secondly, the learned parameters from all the local models are fed to the federated learning server to generate a global federated learning model, which is shared with each client. This cycle is repeated many times to improve threat detection. Federated interpretability weights are computed by aggregating the local interpretability weights of each client. The interpretability module of each client highlights differences in normal and threat sequences that in turn help in the decision-making process. Additionally, it provides forensic investigators with actionable information to effectively analyze system activity and operation.
@@
Bouter et al. [157] present a system named Protoexplorer for visualizing and interpreting predictions of deepfake video data for forensics analysis. The paper mentions the importance of interpretable ML models in high-stakes settings, especially with the prevalence of complex deep-learning models. Also, the paper introduces prototype-based methods as a promising approach to make deep learning interpretable. Protoexplorer improves the explainability of deepfake video detection models by replacing the prototypes with alternative suggested prototypes. Protoexplorer provides a forensics analyst with meaningful visualizations and intuitive interactions with models. This helps the forensics analyst thoroughly evaluate and explain the models. The main action a forensics analyst can perform is zooming in on a part of the image and observing the contribution of the model to the prediction score. This step is called forensic deepfake analysis where a decision is made if the video is manipulated or not. This decision should be interpreted so that it can be presented in a court of law as a piece of trustworthy evidence. For interpretation, the forensics analyst explores and refines the model and provides an explanation and understanding of how the model aided in making the decision.
Alam et al. [206] present SIFT, a machine-learning–based system for automated file-type identification designed for digital forensic analysis. Rather than depending only on traditional signature or header inspection, SIFT employs AI-driven classification to determine file types even when metadata is missing, manipulated, or deliberately concealed. A major contribution is the use of XAI techniques to make the system’s outputs more transparent and defensible in forensic settings. Their approach involves training deep learning models on byte-level features and applying explanation tools such as LIME and SHAP to highlight which regions or attributes influenced each prediction. Khalid et al. [203] introduce a comprehensive framework that integrates XAI across the CF pipeline, from evidence acquisition and feature extraction to model inference, explanation generation, and presentation to investigators or courts. The aim is to standardize processes and enhance the legal and operational defensibility of AI-generated results. Hermosill et al. [207] analyze SHAP and LIME explanations applied to two classifiers, XGBoost and TabNet, for network intrusion detection using the UNSW-NB15 dataset. Their evaluation covers explanation stability, global interpretability, and trade-offs between the two post-hoc methods. They also compare classifier performance, noting that XGBoost achieved higher validation accuracy than TabNet. Khalid et al. [208] investigate combining unsupervised learning with XAI approaches for forensic scenarios characterized by massive unlabeled datasets (e.g., memory forensics, large disk images). Their work emphasizes clustering and anomaly detection augmented with feature-level and prototype-based explanations to reveal novel artifacts and assist investigators in prioritizing analysis.
@@ 6.2.1 Summary
This compilation of research emphasizes the swiftly evolving role of XAI across diverse CF tasks. These tasks encompass log analysis, intrusion detection, deepfake investigations, file-type identification, IoT forensics, and federated systems. The key findings from these studies can be categorized into six primary themes, highlighting the notable impact and advancements that XAI brings to the field of CF.
  1. 1.
    XAI is rapidly becoming a foundational requirement for cyber forensics, driven by legal, ethical, and operational constraints.
  2. 2.
    AI paired with XAI significantly improves forensic workflows, from log analysis and file-type identification to deepfake detection and IoT event reconstruction.
  3. 3.
    Explanations, not just predictions, are crucial for reliability, validation, and courtroom defensibility.
  4. 4.
    Unsupervised and adversarial settings benefit especially from XAI, where explanations reveal hidden patterns, anomalies, and system vulnerabilities.
  5. 5.
    A shift is underway from isolated XAI tools to integrated, pipeline-level frameworks that support end-to-end forensic analysis.
  6. 6.
    Human expertise remains essential, with XAI functioning as an augmentation tool rather than a replacement.
@@ 7 Challenges and Future Works
In this section, we discuss some of the challenges that need to be resolved for the promising development and use of a practical XAI-CF system. We also provide and recommend solutions to these challenges as part of future work. Finally, we introduce a five-layer framework for an XAI-CF system, while leaving the detailed ontology development for future work.
7.1 Human Computer Interaction (HCI)
@@
Interaction with an XAI-CF system allows the user to ask follow-up questions/queries to better understand and evaluate the explanations. Another major reason for interacting with an XAI-CF system is to include a human-in-the-loop that can produce valuable, practical, and structured explanations. To build such systems that are successful in practice, we need to provide the user with a usable interactive XAI-CF system. Human-Computer Interaction (HCI) and XAI are upcoming research areas and very limited research [202, 198] has been reported in this area, and very few [157] on HCI and XAI-CF. As part of the CF research community, this is our responsibility to build the new XAI-CF systems interactive from the ground up. There are already several research works on interacting with AI systems [214, 215, 216, 217, 218]. Recently developed ChatGPT [12, 8], the new AI language model has revolutionized the approach in AI to human-model-interaction. ChatGPT is being used in various domains, such as business, education, and healthcare etc. The potentials of ChatGPT are unlimited, but it can be a double-edged sword [8]. For example, for businesses, it can create creative content such as ideas for advertisements, and on the other hand, it can create hallucinations and produce fake information. One of the future works is to explore the use of ChatGPT to improve HCI in XAI-CF systems and identify fake or false information for its successful use. Consideration of ethical standards and security measures should be paramount in the development process of interactive XAI-CF systems. Francisco Herrera [219] argues that the evolution of XAI is not just a technical problem (making black‐box models interpretable), but fundamentally a sociotechnical challenge. From an HCI standpoint, it’s not enough that a model can explain itself; these explanations must be meaningfully interpretable and actionable by human collaborators. A practical XAI-CF system is human-centered, because without human intervention and understanding CF produces no results. Therefore, HCI for such a system should be guided by human values and needs. One such challenge is socio-technical and is about interacting with an XAI-CF system by a person (e.g., law enforcement community) who has little knowledge of such systems. For example, retraining a model on new data is an advanced skill and requires setting up hyperparameters that influence the learning process. One solution is to transfer such knowledge to the law enforcement community or develop new HCI tools that facilitate and automate such actions with ease and without requiring advanced skills or knowledge. We recommend the latter option, which makes the solution general and is human-centered.
7.2 Bias Management